Outsourcing (Make / Buy)

Kites unlimited uses 50,000 plastic inserts annually to complete their kites. THey can be purchased externally for $0.25. Current per unit production costs for the inserts are:

Direct Material $ 0.03

Direct labor 0.05

Variable Overhead 0.15

Fixed Overhead 10,000.00   

a) Should the inserts be made internally or purchased externally? What is the difference in net income? (Note which is more favorable.)

b) Suppose that one-half of the fixed costs incurred by Kites unlimited could be eliminated if the company outsources the inserts. Should the inserts be made internally or purchased externally? What is the difference in net income? (Note which is more favorable.)
Answer #1

Differential analysis :

Make Buy
Direct material 1500
Direct labour 2500
Variable overhead 7500
Purchase @0.25 12500
Total 11500 12500

Company should make internally difference in net income = (12500-11500) = $1000

b) Differential analysis :

Make Buy
Direct material 1500
Direct labour 2500
Variable overhead 7500
Fixed overhead (10000/2) 5000
Purchase @0.25 12500
Total 16500 12500

Company should buy.  difference in net income = (16500-12500) = $4000